# Complex Valued Deep Operator Network (DeepONet) $[\mathcal{G}]$ for Three Dimensional Maxwell’s Equations: $\mathcal{G}\in\mathbb{C}^{m\times n}$

### v1

@@ Abstract
The Maxwell’s equations, a system of linear partial differential equations (PDEs), describe the behavior of electric and magnetic fields in time and space and is essential for many important electromagnetic applications. Although numerical methods have been applied successfully in the past, the primary challenge in solving these equations arises from the frequency of electromagnetic fields, which depends on the shape and size of the objects to be resolved. The frequency dictates the grid size used for the spatial discretization of the PDEs. Since the domain of influence for these equations is compactly supported, even a small perturbation in frequency necessitates a new discretization of Maxwell’s equations, resulting in substantial computational costs. In this work, we investigate the potential of neural operators, particularly the Deep Operator Network (DeepONet) and its variants, as a surrogate model for Maxwell’s equations. Existing DeepONet implementations are restricted to real-valued data in $\mathbb{R}^{n}$, but since the time-harmonic Maxwell’s equations yield solutions in the complex domain $\mathbb{C}^{n}$, a specialized architecture is required to handle complex algebra. We propose a formulation of DeepONet for complex data, define the forward pass in the complex domain, and adopt a reparametrized version of DeepONet for more efficient training. We also propose a unified framework to combine a plurality of DeepONets, trained for multiple electromagnetic field components, to incorporate the boundary condition. As shown by computational experiments, our method significantly enhances the efficiency of predicting scattered fields from a spherical object at arbitrary high frequencies.
Keywords Maxwell’s equations  $\cdot$ Neural operators  $\cdot$ Complex-valued neural network  $\cdot$ Radar cross section
@@ 1 Introduction:
Electromagnetics is a fundamental subject that studies electric and magnetic fields, their interactions, and their effects on materials and systems, which is crucial to a variety of daily and industrial applications, including wireless communication [1, 2, 3], radar [4, 5, 6], and medical imaging [7, 8]. Central to this field are Maxwell’s equations, which describe how electromagnetic waves propagate and respond to external charges and currents. However, analytical solutions in closed form are known for only a very limited number of special and trivial cases. This limitation has motivated the development of various high order numerical methods into solve the Maxwell’s equation in time and frequency domain [9, 1, 10]. However, because the grid resolution depends on the size of the object being resolved, the numerical methods can become computationally demanding due to the high frequencies and small skin depth required to resolve the small objects. To address this issue, we propose using a neural network-based approximate operator as a surrogate for the Maxwell’s equation solver.
@@
Recently, computational problems in science and engineering have seen significant advancements through the application of deep learning methods. Commonly used neural network solvers can be broadly divided into two types: the first type focuses on learning mappings between input $\mathbb{R}^{m}$ and output data ${\mathbb{R}^{m}}$, while the second type, known as neural operators, learns mapping between two functional spaces. In this work, we focus on the latter—neural operators [11, 12, 13]—which aim to approximate unknown operators that are often linked to the solution operator to differential equations. In mathematical terms, given data pairs $(v,u)$, where $v\in\mathcal{V}$ and $u\in\mathcal{U}$ belong to two function spaces, and a potentially nonlinear operator $\mathcal{G}:\mathcal{V}\mapsto\mathcal{U}$ such that $\mathcal{G}(v)=u$, the objective is to find a neural network approximation of $\mathcal{G}$, denoted by $\hat{\mathcal{G}}$, such that it can generalize to unseen data. By leveraging neural operators, we can efficiently replace highly complex and computationally intensive multiphysics systems and deliver functional outputs in real-time.
@@
Deep Operator Network (DeepONet) [12] is a neural network based parametrized model for learning nonlinear operators and capturing the inherent relationships between input and output functions. DeepONets are built upon the universal operator approximation theorem [14, 12], which is an infinite dimensional analogue of the universal approximation theorem for neural networks [15, 16].An architectural diagram of a DeepONet with the commonly used nomenclature for its components is shown in Figure 1. The DeepONet is a two-pronged deep learning network consisting of a branch network, which can take a multi-fidelity or multi-modal input [17, 18, 19, 20], and a trunk network encoded the independent variables defining the output space, e.g., in space–time coordinates. To recover the solution of a PDE the output of branch network is project on the output of trunk networks via an inner-product. DeepONet has seen success in solving complex problems in diverse scientific and engineering problems [21, 22, 23, 24, 25, 26], climate [27, 28, 29], and material science [30, 31, 18]. In Figure 1, we have shown fully connected architecture but any architecture can be used depending on the dimensionality of the training and testing data.
@@
Figure 1: A schematic representation of a DeepONet, which is a neural operator trained to learn the mapping from the input function $u(x)$ to the output function $\mathcal{G}(u)(y)$, evaluated at $y\in\mathbb{R}^{q}$. DeepONet consists of two networks: a branch network, which processes the input function evaluations $[u(x_{1}),u(x_{2}),\dots,u(x_{m})]^{\top}$, and a trunk network, which takes in the coordinates $y$.
@@
All the existing applications of DeepONets focus on real-valued input and output data. However, many scientific problems involve complex-valued data, particularly in the solutions of important partial differential equations such as the Schrodinger equation [32, 33], Helmholtz equation [34], and Maxwell’s equations [35, 36]. Although, there are studies of solving Maxwells equation in 2D using PINNs [37, 38, 39, 40] but they consider the system with real valued solutions. To address the challenge of representing both the magnitude ($m$) and phase ($\varphi$) of a complex number $z=me^{i\varphi}\in\mathbb{C}$ and take advantage of the complex algebra, this work introduces the complex-valued DeepONet for learning operator mappings between complex-valued functions as input and output. Specifically, we demonstrate the effectiveness of using the complex-valued surrogate model for electromagnetic computations governed by the system of time-harmonic 3D Maxwell’s equations, where the DeepONet is leveraged to learn the general mapping from incident fields to the scattered fields parametrized by an arbitrary angular frequency.
@@
To incorporate the the complex valued input and output in a neural network requires redefining all the three steps (1) Affine Transform (2) Non-linearity and (3) backpropagation. The first step requires matrix-vector product, which is well defined for complex valued matrices. The second step will require defining the non-linear activation function $f:\mathbb{C}^{n}\to\mathbb{C}^{m}$ for complex valued variables as domain and co-domain. This will require defining the new complex valued activation function. However, there are existing definitions of such functions and in this study we use the definition proposed by [41]. In third step we require computing the gradient of loss function with respect to complex parameters of neural network. To achieve this efficiently, we have used Wirtinger calculus [42]. The usage of complex valued neural network is not common, but a few studies have shown great promise of using complex valued model for classification [43] and regression [44, 45, 46] problems.
@@
This paper is structured as follows: In section 2, we present the definition of the problem, governing equations, the geometry of the domain, and the data generation techniques used to train n the model. Next, in section 3, we describe the architecture and forward passes for the complex-valued DeepONet. In addition, we adapts a two-step training approach to reparametrize the complex-valued network and demonstrate how separate DeepONets can be combined to enforce the boundary conditions for a perfect electric conductor (PEC). Lastly, in section 4, we highlight the superior results of the complex-valued model compared to its real-valued counterpart, along with the post-processed radar cross-section (RCS) results.
@@ 2.1 System of time-harmonic Maxwell’s equations
×$\nabla\times\mu^{-1}\nabla\times\boldsymbol{E}-\kappa^{2}\varepsilon\boldsymbol{E}=0\quad\text{in }\Omega,$ (3)
@@
$\boldsymbol{v}\times\mu^{-1}\nabla\times\boldsymbol{E}-ikZ\boldsymbol{E}_{T}$ $=Q\big{(}\boldsymbol{v}\times\mu^{-1}\nabla\times\boldsymbol{E}+ikZ\boldsymbol{E}_{T}\big{)}+\boldsymbol{g}\quad\text{on }\,\partial\Omega,$ (4)
where, $\boldsymbol{E}_{T}=\boldsymbol{v}\times(\boldsymbol{E}\times\boldsymbol{v})$ is the tangential component of the electric field, $Q$ is a complex scalar-valued function on the boundary with $\lvert Q\rvert\leq 1$, and $\boldsymbol{g}$ is a prescribed source function. The $Z(\geq 0)$ is defined on the boundary $\partial\Omega$, and we take it to be $Z=\sqrt{\lvert\mu\rvert/\lvert\varepsilon\rvert}$. Note that the choice $Q=1$ corresponds to a Perfect Electric Conductor (PEC) type boundary condition for the waves scattered of the PEC surface and recovered as
@@
For scattering applications, we assume that we have a bounded scatterer which can be impenetrable (for example, with an impedance or PEC boundary condition) or penetrable (for example, a dielectric) occupying a domain $D$ in the interior of the computational domain $\Omega$. A known incident field $\boldsymbol{E}^{i}$ impinges on the scatterer and creates an outgoing scattered field $\boldsymbol{E}^{s}$. Then, the total field $\boldsymbol{E}=\boldsymbol{E}^{i}+\boldsymbol{E}^{s}$ satisfies the homogeneous Maxwell system 1. The incident field $\boldsymbol{E}^{i}$ is assumed to satisfy the background Maxwell’s equations in the neighborhood of $D$, which is expressed as
××$\nabla\times\nabla\times\boldsymbol{E}^{i}-k^{2}\boldsymbol{E}^{i}=0.$ (6)
In this work, we consider the incident plane wave $\boldsymbol{E}^{i}=\boldsymbol{E}_{0}^{i}\exp(ik\boldsymbol{d}\cdot\boldsymbol{x})$, where the real direction vector $\boldsymbol{d}$ satisfies $\|\boldsymbol{d}\|=1$ and the real polarization $\boldsymbol{E}_{0}^{i}\neq 0$ is such that $\boldsymbol{E}_{0}^{i}\cdot\boldsymbol{d}=0$.
@@
2.2 Computational domain and data generation
[…]
$\lambda=\frac{c}{f}$ (7)
[…]
The training dataset contains the incident fields $\boldsymbol{E}^{i}(x,y,z;f)$ impinging the sphere at different incidence angles, frequencies $(f)$, and spatial locations of sensors, which in this case are coordinates of vertices $(x,y,z)$ of the tetrahedral and triangles. The output of the DeepONet is the total field $\boldsymbol{E}(x,y,z;f)$. To generate the training and testing data, we utilized the Ultra Weak Variational Formulation (UWVF) method [47, 48, 49] to solve the system described by Equation 1. UWVF is a special case of a Trefftz-discontinuous Galerkin (Trefftz-DG) method provided the scattering medium is lossless. The UWVF approach uses an unstructured finite element computational grid and a superposition of plane wave solutions of Maxwell’s equations on each tetrahedral element [50]. UWVF has been found to be efficient for approximating solutions of Maxwell’s equations for a wide range of test problems as shown by Lahivaara et al.[51].
(b) Meshed geometry of the computational domain.
@@ 3.1 Formulation of a real valued DeepONet
where $b_{i}$ and $t_{i}$ are the $r$-dimensional outputs of the branch and trunk networks, respectively. $b_{0}\in\mathbb{R}$ is a trainable bias term. By evaluating the network’s output at the coordinates $\{y_{1},y_{2},\ldots,y_{q}\}$, we obtain the continuous prediction of $\bm{v}$ projected at $\bm{y}$ as $[v(y_{1}),v(y_{2}),\ldots,v(y_{q})]^{\top}$, which approximated the ground truth data.
@@ 3.2 Formulation of a Complex-valued DeepONet
and then recombines the real part $a_{j}$ and imaginary part $b_{j}$ to produce the complex-valued output $a_{j}+ib_{j}$. However, this separate representation does not respect the phase information that is captured by complex algebra. In contrast, a complex-valued neural network directly learns the complex-valued mapping
@@
Previous research [43, 44, 45, 46] has shown that the real-valued approach to inherently complex-valued data performs poorly in terms of efficient architecture, convergence, and generalization ability. Since the solutions to the time-harmonic Maxwell’s equations are inherently complex-valued, we address this challenge by implementing the DeepONet in the complex domain to learn the mapping between a complex-valued fields. The complex-valued DeepONet retains the same structure as the basic DeepONet, but both the input and output functions are complex-valued, and all calculations are performed using the complex64 datatype.
@@
periodic feature of $\text{periodic feature of }\boldsymbol{y}=\boldsymbol{y}+\sum_{n=1}^{N}\sin(n\boldsymbol{y})+\cos(n\boldsymbol{y}),$ (13)
@@
The same MSE function is used for all other components of the electric and magnetic fields. A separate DeepONet is implemented for each of the six components of the electric and magnetic fields to minimize their losses separately. Our implementation utilizes the Complex Valued Neural Network package [53] to simplify the workflow.

### v2

# Complex Valued Deep Operator Network (DeepONet) $\mathcal{G}]$ for Three Dimensional Maxwell’s Equations: $\mathcal{G}\in\mathbb{C}^{m\times n}$

@@ Abstract
Maxwell’s equations, a system of linear partial differential equations (PDEs), describe the behavior of electric and magnetic fields in time and space and are essential for many important electromagnetic applications. Although numerical methods have been applied successfully in the past, the primary challenge in solving these equations arises from the frequency of electromagnetic fields, which depends on the shape and size of the objects to be resolved. The frequency dictates the grid size used for the spatial discretization of the PDEs. Since the domain of influence for these equations is compactly supported, even a small perturbation in frequency necessitates a new discretization of Maxwell’s equations, resulting in substantial computational costs. In this work, we investigate the potential of neural operators, particularly the Deep Operator Network (DeepONet) and its variants, as a surrogate model for Maxwell’s equations. Existing DeepONet implementations are restricted to real-valued data in $\mathbb{R}^{n}$, but since the time-harmonic Maxwell’s equations yield solutions in the complex domain $\mathbb{C}^{n}$, a specialized architecture is required to handle complex algebra. We propose a formulation of DeepONet for complex data, define the forward pass in the complex domain, and adopt a reparametrized version of DeepONet for more efficient training. We also propose a unified framework to combine a plurality of DeepONets, trained for multiple electromagnetic field components, to incorporate the boundary condition. We conduct computational experiments on a three-dimensional metallic sphere without singularities and on a metallic almond-shaped target to demonstrate the effectiveness of the proposed method for problems involving singularity-prone solutions. As shown by computational experiments, our method significantly enhances the efficiency of predicting scattered fields from a spherical object at arbitrary high frequencies.
Keywords Maxwell’s equations $\cdot$ Radar cross section $\cdot$ Physics-informed machine learning $\cdot$ Neural operators $\cdot$ Complex-valued neural network
@@ 1 Introduction:
Electromagnetics is a fundamental subject that studies electric and magnetic fields, their interactions, and their effects on materials and systems, which is crucial to a variety of daily and industrial applications, including wireless communication [7, 43, 47], radar [56, 14, 6], and medical imaging [50, 1]. Central to this field are Maxwell’s equations, which describe how electromagnetic waves propagate and respond to external charges and currents. However, analytical solutions in closed form are known for only a very limited number of special and trivial cases. This limitation has motivated the development of various high order numerical methods to solve the Maxwell’s equation in time and frequency domain [21, 7, 41]. However, because the grid resolution depends on the size of the object being resolved, the numerical methods can become computationally demanding due to the high frequencies and small skin depth required to resolve the small objects. To address this issue, we propose using a neural network-based approximate operator as a surrogate for the Maxwell’s equation solver.
@@
Recently, computational problems in science and engineering have seen significant advancements through the application of deep learning methods. Commonly used neural network solvers can be broadly divided into two types: the first type focuses on learning mappings between input $\mathbb{R}^{m}$ and output data $\mathbb{R}^{m}$}, while the second type, known as neural operators, learns mappings between two functional spaces. In this work, we focus on the latter—neural operators [29, 38, 8]—which aim to approximate unknown operators that are often linked to the solution operator to differential equations. In mathematical terms, given data pairs $(v,u)$, where $v\in\mathcal{V}$ and $u\in\mathcal{U}$ belong to two function spaces, and a potentially nonlinear operator $\mathcal{G}:\mathcal{V}\mapsto\mathcal{U}$ such that $\mathcal{G}(v)=u$, the objective is to find a neural network approximation of $\mathcal{G}$, denoted by $\hat{\mathcal{G}}$, such that it can generalize to unseen data. By leveraging neural operators, we can efficiently replace highly complex and computationally intensive multiphysics systems and deliver functional outputs in real-time.
@@
Deep Operator Network (DeepONet) [38] is a neural network based parametrized model for learning nonlinear operators and capturing the inherent relationships between input and output functions. DeepONets are built upon the universal operator approximation theorem [10, 38], which is an infinite dimensional analogue of the universal approximation theorem for neural networks [12, 22]. An architectural diagram of a DeepONet with the commonly used nomenclature for its components is shown in Figure 1. The DeepONet is a two-pronged deep learning network consisting of a branch network, which can take a multi-fidelity or multi-modal input [13, 39, 23, 55], and a trunk network that encodes the independent variables defining the output space, e.g., in space–time coordinates. To recover the solution of a PDE the output of branch network is projected on the output of trunk networks via an inner-product. DeepONet has seen success in solving complex problems in diverse scientific and engineering problems [34, 16, 48, 9, 18, 49], climate [28, 20, 52], and material science [17, 19, 39]. In Figure 1, we have shown fully connected architecture but any architecture can be used depending on the dimensionality of the training and testing data.
@@
Figure 1: A schematic representation of a DeepONet, which is a neural operator trained to learn the mapping from the input function $u(x)$ to the output function $\mathcal{G}(u)(y)$, evaluated at $y\in\mathbb{R}^{q}$. DeepONet consists of two networks: a branch network, which processes the input function evaluations $u(x_{1}),u(x_{2}),\dots,u(x_{m})]^{\top}$, and a trunk network, which takes in the coordinates $y$.
@@
All the existing applications of DeepONets focus on real-valued input and output data. However, many scientific problems involve complex-valued data, particularly in the solutions of important partial differential equations such as the Schrodinger equation [5, 57], Helmholtz equation [15], and Maxwell’s equations [42, 46]. Although there are studies of solving Maxwells equation in 2D using PINNs [11, 53, 54, 33], they consider the system with real valued solutions. To address the challenge of representing both the magnitude ($m$) and phase ($\varphi$) of a complex number $z=me^{i\varphi}\in\mathbb{C}$ and take advantage of the complex algebra, this work introduces the complex-valued DeepONet for learning operator mappings between complex-valued functions as input and output. Specifically, we demonstrate the effectiveness of using the complex-valued surrogate model for electromagnetic computations governed by the system of time-harmonic 3D Maxwell’s equations, where the DeepONet is leveraged to learn the general mapping from incident fields to the scattered fields parametrized by an arbitrary angular frequency.
@@
To incorporate the complex valued input and output in a neural network requires redefining all the three steps (1) Affine Transform (2) Non-linearity and (3) backpropagation. The first step requires matrix-vector product, which is well defined for complex valued matrices. The second step will require defining the non-linear activation function $f:\mathbb{C}^{n}\to\mathbb{C}^{m}$ for complex valued variables as domain and co-domain. This will require defining the new complex valued activation function. However, there are existing definitions of such functions and in this study we use the definition proposed by [27]. In the third step we require computing the gradient of loss function with respect to complex parameters of neural network. To achieve this efficiently, we have used Wirtinger calculus [2]. The usage of complex valued neural network is not common, but a few studies have shown great promise of using complex valued model for classification [37] and regression [51, 3, 44] problems.
@@
This paper is structured as follows: In section 2, we present the definition of the problem, governing equations, the geometry of the domain, and the data generation techniques used to train the model. Next, in section 3, we describe the architecture and forward passes for the complex-valued DeepONet. In addition, we adapt a two-step training approach to reparametrize the complex-valued network and demonstrate how separate DeepONets can be combined to enforce the boundary conditions for a perfect electric conductor (PEC). Lastly, in section 4, we highlight the superior results of the complex-valued model compared to its real-valued counterpart, along with the post-processed radar cross-section (RCS) results.
1.1 Contributions:
  1. 1.
    In this work, we propose a composite neural network framework in which learning is carried out directly in the complex space $\mathbb{C}^{N}$. Unlike traditional approaches that split complex-valued solutions into their real and imaginary components and process them using separate real-valued neural networks, our method avoids such decomposition, which can obscure the inherent coupling between the two components and introduce unnecessary architectural complexity. Instead, we operate natively in the complex domain by employing complex-valued activation functions and developing a corresponding backpropagation algorithm defined entirely in complex space. This unified formulation preserves the intrinsic structure of complex-valued representations and enables more faithful and efficient learning for problems where the governing physics is naturally expressed in complex form.
  2. 2.
    We train six separate DeepONets to model the electric field ($\bm{E}$) and magnetic field ($\bm{H}$) components individually, with the goal of achieving improved representation and approximation accuracy for each component. However, in order to properly impose the underlying physics on the DeepONet predictions, it is necessary to consider all six components simultaneously. To address this challenge, we propose a novel approach that combines these six individual DeepONets into a unified framework, allowing us to enforce the physics of electromagnetic waves through the governing partial differential equations (PDEs). This strategy ensures that the DeepONet predictions remain physically consistent while benefiting from the specialized learning of each field component.
  3. 3.
    In this study, we also demonstrate the applicability of the proposed DeepONet to a three-dimensional almond-shaped complex geometry that gives rise to singularities. Furthermore, we address the training of DeepONets for complex geometries, such as the almond shape with two singularity locations, highlighting the model’s capability to handle multiple singularities simultaneously.
@@ 2.1 System of time-harmonic Maxwell’s equations
$\nabla\times\mu^{-1}\nabla\times\boldsymbol{E}-\kappa^{2}\varepsilon\boldsymbol{E}=0\quad\text{in }\Omega,$ (3)
@@
$\bm{\hat{n}}\times\mu^{-1}\nabla\times\boldsymbol{E}-i\kappa Z\boldsymbol{E}_{T}$ $=Q\big(\bm{\hat{n}}\times\mu^{-1}\nabla\times\boldsymbol{E}+i\kappa Z\boldsymbol{E}_{T}\big)+\boldsymbol{g}\quad\text{on }\,\partial\Omega,$ (4)
[…]
@@
For scattering applications, we assume that we have a bounded scatterer which can be impenetrable (for example, with an impedance or PEC boundary condition) or penetrable (for example, a dielectric) occupying a domain $D$ in the interior of the computational domain $\Omega$. A known incident field $\boldsymbol{E}^{i}$ impinges on the scatterer and creates an outgoing scattered field $\boldsymbol{E}^{s}$. Then, the total field $\boldsymbol{E}=\boldsymbol{E}^{i}+\boldsymbol{E}^{s}$ satisfies the homogeneous Maxwell system (1). The incident field $\boldsymbol{E}^{i}$ is assumed to satisfy the background Maxwell’s equations in the neighborhood of $D$, which is expressed as
$\nabla\times\nabla\times\boldsymbol{E}^{i}-\kappa^{2}\boldsymbol{E}^{i}=0.$ (6)
In this work, we consider the incident plane wave $\boldsymbol{E}^{i}=\boldsymbol{E}_{0}^{i}\exp(i\kappa\boldsymbol{d}\cdot\boldsymbol{x})$, where the real direction vector $\boldsymbol{d}$ satisfies $\|\boldsymbol{d}\|=1$ and the real polarization $\boldsymbol{E}_{0}^{i}\neq 0$ is such that $\boldsymbol{E}_{0}^{i}\cdot\boldsymbol{d}=0$.
@@ 3.1 Formulation of a real valued DeepONet
where $b_{i}$ and $t_{i}$ are the $r$-dimensional outputs of the branch and trunk networks, respectively. $b_{0}\in\mathbb{R}$ is a trainable bias term. By evaluating the network’s output at the coordinates $\{y_{1},y_{2},\ldots,y_{q}\}$, we obtain the continuous prediction of $\bm{v}$ projected at $\bm{y}$ as $v(y_{1}),v(y_{2}),\ldots,v(y_{q})]^{\top}$, which approximates the ground truth data.
@@ 3.2 Formulation of a Complex-valued DeepONet
and then recombines the real part $a_{j}$ and imaginary part $b_{j}$ to produce the complex-valued output $a_{j}+ib_{j}$. This separate representation may not capture the phase and magnitude relationships as effectively as treating the complex number as a unified entity (see Appendix A for an illustrative example). In contrast, a complex-valued neural network directly learns the complex-valued mapping
@@
Previous research [37, 51, 3, 44] has shown that the real-valued approach to inherently complex-valued data performs poorly in terms of efficient architecture, convergence, and generalization ability. Since the solutions to the time-harmonic Maxwell’s equations are inherently complex-valued, we address this challenge by implementing the DeepONet in the complex domain to learn the mapping between complex-valued fields. The complex-valued DeepONet retains the same structure as the basic DeepONet, but both the input and output functions are complex-valued, and all calculations are performed using the complex64 datatype.
@@
$\text{periodic feature of }\boldsymbol{y}=\boldsymbol{y}+\sum_{n=1}^{N}\sin(n\boldsymbol{y})+\cos(n\boldsymbol{y}),$ (12)
@@
The same MSE function is used for all other components of the electric and magnetic fields. A separate DeepONet is implemented for each of the six components of the electric and magnetic fields to minimize their losses separately. Our implementation utilizes the Complex Valued Neural Network package [4] to simplify the workflow. To illustrate the advantages of the complex-valued DeepONet over the real-valued counterpart, we present a detailed analysis along with computational results for a representative complex-domain function approximation. The complete comparison and supporting results are provided in Appendix A, highlighting the improved accuracy and stability achieved by the complex-valued architecture.